OpenStudy (kennnnziee):

A 6 inch personal pizza has 600 calories, with 240 of those from fat. A 16 inch pizza is cut into 8 slices. Estimate the number of calories in one slice of a 16 inch pizza.

OpenStudy (peachpi):

I think you may need to scale this by area. \[\frac{ 6"area }{ 16"area }=\frac{ 6"calories }{ 16"calories }\] So for calories you'd have \[\frac{ \pi(6^2) }{ \pi(16^2) }=\frac{ 600 }{ x }\] or\[\frac{ 36 }{ 256 }=\frac{ 600 }{ x }\] x is the calories in a whole 16" pizza, so divide by 8 to get the calories in one slice. Then do the same for fat calories, except change 600 to 240

OpenStudy (kennnnziee):

thank you so much! i understand it now!